# Jacobian Descent For Multi-Objective Optimization

### v1

@@ 4 Existing aggregators
In the context of multi-task learning, several works have proposed iterative optimization algorithms based on the combination of task-specific gradients [10, 11, 13, 12, 14, 15, 16]. Aggregators can be extracted from these methods, thereby formulating them as variants of JD. More specifically, since the gradients are stochastically estimated from batches of data, these are cases of what we call SEJD. In the following, we briefly present the most prominent aggregators and summarize their properties in Table 1. We also consider $\mathcal{A}_{\text{Mean}}$, which simply averages the rows of the Jacobian. Their formal definitions are provided in Appendix B. Some of them are also illustrated in Figure 1(b).
@@
$\mathcal{A}_{\text{RGW}}$ aggregates the matrix using a random vector of weights [14]. $\mathcal{A}_{\text{MGDA}}$ gives the aggregation that maximizes the smallest improvement [6, 10, 5]. $\mathcal{A}_{\text{CAGrad}}$ maximizes the smallest improvement in a ball around the average gradient whose radius is parameterized by $c\in\left[0,1\right[{}$ [12]. $\mathcal{A}_{\text{PCGrad}}$ projects each gradient onto the orthogonal hyperplane of other gradients in case of conflict, iteratively and in a random order [11]. It is, however, only non-conflicting when $m\varleq 2$, in which case $\mathcal{A}_{\text{PCGrad}}=m\cdot\mathcal{A}_{\text{UPGrad}}$. IMTL-G is a method to balance some gradients with impartiality [13]. It is only defined for linearly independent gradients, but we generalize it as a formal aggregator, denoted $\mathcal{A}_{\text{IMTL-G}}$, in Appendix B.6. Aligned-MTL orthonormalizes the Jacobian and weights its rows according to some preferences [16]. We denote by $\mathcal{A}_{\text{Aligned-MTL}}$ this method with uniform preferences. $\mathcal{A}_{\text{Nash-MTL}}$ aggregates Jacobians by finding the Nash equilibrium between task-specific gradients [15]. Lastly, the GradDrop layer [17] defines a custom backward pass that combines gradients with respect to some internal activation. The corresponding aggregator, denoted $\mathcal{A}_{\text{GradDrop}}$, randomly drops out some gradient coordinates based on their sign and sums the remaining ones.
@@
In the context of continual learning, to limit forgetting, an idea is to project the gradient onto the dual cone of gradients computed with past examples [33]. This idea can be translated into an aggregator that projects the mean gradient onto the dual cone of the rows of the Jacobian. We name this $\mathcal{A}_{\text{DualProj}}$.
@@
Several other works consider the gradients to be noisy when making their theoretical analysis [34, 35, 36, 37, 38]. Their solutions for combining gradients are typically stateful. Unifying their formulations would thus require a more complex variant of Jacobian descent.
@@
In the federated learning setting, several aggregators have been proposed to combine the model updates while being robust to adversaries [39, 40, 41, 42]. We do not study them here as their desired properties mainly revolve around security.
@@
[6, 10, 5] $\mathcal{A}_{\text{MGDA}}$
[33] $\mathcal{A}_{\text{DualProj}}$
[11] $\mathcal{A}_{\text{PCGrad}}$
[17] $\mathcal{A}_{\text{GradDrop}}$
[13] $\mathcal{A}_{\text{IMTL-G}}$
[12] $\mathcal{A}_{\text{CAGrad}}$
[14] $\mathcal{A}_{\text{RGW}}$
[15] $\mathcal{A}_{\text{Nash-MTL}}$
[16] $\mathcal{A}_{\text{Aligned-MTL}}$
@@ 5 Experiments
In the following, we present empirical results for instance-wise risk minimization on some simple image classification datasets. IWRM is performed by stochastic sub-Jacobian descent, as described in Section 3. A key consideration is that when the aggregator is $\mathcal{A}_{\text{Mean}}$, this approach becomes equivalent to empirical risk minimization with SGD. Thus, it constitutes a baseline for comparison.
We train convolutional neural networks on subsets of SVHN [43], CIFAR-10 [44], EuroSAT [45], MNIST [46], Fashion-MNIST [47] and Kuzushiji-MNIST [48]. To make the comparisons as fair as possible, we have tuned the learning rate very precisely for each aggregator, as explained in detail in Appendix C.1. To keep computational costs reasonable, we have thus limited the size of each training dataset to 1024 images. This also enables re-running the same experiments several times on different subsets and with different seeds to gain confidence in our results. Note that this is strictly an optimization problem: we are not studying the generalization of the model, which would be captured by some performance metric on a test set. Other experimental settings, such as the network architectures and the total computational budget used to run our experiments, are given in Appendix C. Figure 2 reports the main results on SVHN and CIFAR-10, two of the datasets exhibiting the most substantial performance gap. Results on the other datasets and aggregators are reported in Appendix D.1. They also demonstrate a significant performance gap.
@@
Figure 2: Optimization metrics obtained with IWRM with 1024 training examples and a batch size of 32, averaged over 8 random runs. The shaded area around each curve shows the estimated standard error of the mean over the 8 runs. Curves are smoothed for readability. Best viewed in color.
Here, we compare the aggregators in terms of their average loss over the training set. This is precisely the goal of ERM. For this reason, it is quite surprising to see that the direct optimization of this objective, as performed by $\mathcal{A}_{\text{Mean}}$, can be outperformed by some other aggregators. In particular, $\mathcal{A}_{\text{UPGrad}}$, and to some extent $\mathcal{A}_{\text{DualProj}}$, provide improvements on all datasets. Figures 2(b) and  2(d) show the similarity between the update of each aggregator and the update given by $\mathcal{A}_{\text{Mean}}$. As explained in Section 2.3, if there is no conflict among gradients, $\mathcal{A}_{\text{UPGrad}}$ becomes equivalent to $\mathcal{A}_{\text{Mean}}$. The similarity curve of $\mathcal{A}_{\text{UPGrad}}$ thus suggests substantial conflict between the individual gradients, especially early into the training. Our interpretation is that at the beginning, the gradients of hard examples are dominated by those of easier examples if averaged. However, with $\mathcal{A}_{\text{UPGrad}}$, the updates will positively affect hard examples even during the early training phase. Since fitting such examples is more complex and time-consuming, it is beneficial to consider them early on. Another possibility is that the projections of $\mathcal{A}_{\text{UPGrad}}$ improve its stability compared to $\mathcal{A}_{\text{Mean}}$, making it favor a learning rate giving larger updates, thus enabling a faster convergence. The sub-optimal performance of $\mathcal{A}_{\text{MGDA}}$ in this setting can be attributed to its sensitivity to small gradients. If any row of the Jacobian approaches zero, the aggregation by $\mathcal{A}_{\text{MGDA}}$ will also approach zero. Note that an advantage of linearity under scaling is to explicitly prevent this from happening.
Overall, these experiments demonstrate a high potential for the IWRM paradigm and confirm the validity of JD, and more specifically of SSJD, as multi-objective optimization algorithms. More generally, they question the soundness of the ubiquitous loss scalarization in deep learning, especially when conflict is significant. Besides, the superiority of $\mathcal{A}_{\text{UPGrad}}$ in such a simple setting confirms our theoretical findings on the desired properties of aggregators. Consequently, using $\mathcal{A}_{\text{UPGrad}}$ would likely be beneficial as an out-of-the-box replacement whenever some gradients are averaged or summed. An interesting consideration is that with SGD, increasing the batch size decreases the variance, but with SSJD combined with $\mathcal{A}_{\text{UPGrad}}$, the effect is non-trivial since it also narrows the dual cone. Additional results obtained when varying the batch size or when updating the parameters with the Adam optimizer are available in Appendices D.2 and D.3, respectively.
It is important to note that an iteration of SSJD takes more time than an iteration of SGD. The exact runtime of SSJD depends on many factors, such as the aggregator, the parallelization ability of the device on which the Jacobians are computed, and the implementation. We provide an empirical comparison of the speed of SGD and SSJD with different aggregators in Appendix E. Besides, we address computational efficiency concerns in the next section.
6 Towards an efficient implementation
@@
When the number of objectives is dominated by the number of parameters of the model, the main overhead of JD comes from the computation of a Jacobian matrix rather than a single gradient. We give several paths towards faster algorithms.
Grouping objectives
Linear scalarization of a multi-objective optimization problem casts it as single-objective. This idea can be relaxed by considering $k$ different scalarizations of the objectives, with $k<m$ [49]. In the context of multi-task learning, several studies have already worked on task relationships and groupings [50, 51, 52, 53, 54]. We believe combining objectives with generally aligned gradients would be preferable.
Gramian-based JD
@@
In the following, we show how to make a step of JD without even having to compute the Jacobian. For any $J\in\mathbb{R}^{m\times n}$, the matrix $G=JJ^{\top}$ is called the Gramian of $J$ and is positive semi-definite. Let $\mathcal{M}_{m}\subseteq\mathbb{R}^{m\times m}$ be the set of positive semi-definite matrices. The Gramian of the Jacobian, denoted $\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}})=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})^{\top}\in\mathcal{M}_{m}$, captures the relations – including conflicts – between all pairs of gradients.
Whenever $\mathcal{A}$ is a weighted aggregator, the update of JD is ${\boldsymbol{y}}=-\eta\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})^{\top}{\boldsymbol{w}}$ for some vector of weights ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$. Substituting this into the Taylor approximation of (1) gives
${\boldsymbol{f}}({\boldsymbol{x}}+{\boldsymbol{y}})={\boldsymbol{f}}({\boldsymbol{x}})-\eta\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{w}}+o\left(\eta\sqrt{{\boldsymbol{w}}^{\top}\cdot\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{w}}}\right).$ (8)
@@
This expression only depends on the Jacobian through its Gramian. It is thus sensible to focus on aggregators whose weights are only function of the Gramian. Denoting this function as $\mathcal{W}~{}:~{}\mathcal{M}_{m}\to{\mathbb{R}}^{m}$, those aggregators are such that $\mathcal{A}(J)=J^{\top}\cdot\mathcal{W}(G)$. Remarkably, all weighted aggregators of Table 1 can be expressed in this form. For such aggregators, substitution and linearity of differentiation [^3] For any ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ and any ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$, $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})^{\top}{\boldsymbol{w}}=\nabla({\boldsymbol{w}}^{\top}{\boldsymbol{f}})({\boldsymbol{x}})$ then yield
$\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))=\nabla\Big{(}\mathcal{W}(\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}}))^{\top}\cdot{\boldsymbol{f}}\Big{)}({\boldsymbol{x}}).$ (9)
@@
After computing $\mathcal{W}(\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$, a step of JD would thus only require the backpropagation of a scalar function. The computational cost of applying $\mathcal{W}$ depends on the aggregator and is often dominated by the cost of computing the Gramian. We now outline an promising alternative algorithm for the latter.
Similarly to the backpropagation algorithm, the chain rule can be leveraged to propagate the Gramian. Let ${\boldsymbol{g}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{k}$ and ${\boldsymbol{f}}:{\mathbb{R}}^{k}\to{\mathbb{R}}^{m}$, then for any ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$,
$\mathcal{G}\kern-1.0pt({\boldsymbol{f}}\circ{\boldsymbol{g}})({\boldsymbol{x}})=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{g}}({\boldsymbol{x}}))\cdot\mathcal{G}\kern-1.0pt{\boldsymbol{g}}({\boldsymbol{x}})\cdot\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{g}}({\boldsymbol{x}}))^{\top}.$ (10)
@@
Besides, when the function has multiple inputs, the Gramian can be computed as a sum of individual Gramians. Let ${\boldsymbol{f}}:\mathbb{R}^{n_{1}+\cdots+n_{k}}\to{\mathbb{R}}^{m}$ and ${\boldsymbol{x}}=\begin{bmatrix}{\boldsymbol{x}}_{1}^{\top}&\cdots&{\boldsymbol{x}}_{k}^{\top}\end{bmatrix}^{\top}$. We can write $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})$ as the concatenation of Jacobians $\begin{bmatrix}\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}({\boldsymbol{x}})&\cdots&\mathcal{J}\kern-2.0pt_{k}{\boldsymbol{f}}({\boldsymbol{x}})\end{bmatrix}$, where $\mathcal{J}\kern-2.0pt_{i}{\boldsymbol{f}}({\boldsymbol{x}})$ is the Jacobian of ${\boldsymbol{f}}$ with respect to ${\boldsymbol{x}}_{i}$ evaluated at ${\boldsymbol{x}}$. For any $i\in[m]$, let $\mathcal{G}\kern-1.0pt_{i}{\boldsymbol{f}}({\boldsymbol{x}})=\mathcal{J}\kern-2.0pt_{i}{\boldsymbol{f}}({\boldsymbol{x}})\mathcal{J}\kern-2.0pt_{i}{\boldsymbol{f}}({\boldsymbol{x}})^{\top}$. Then
$\mathcal{G}\kern-1.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{1},\dots,{\boldsymbol{x}}_{k})=\sum_{i=1}^{k}\mathcal{G}\kern-1.0pt_{i}{\boldsymbol{f}}({\boldsymbol{x}}_{1},\dots,{\boldsymbol{x}}_{k}).$ (11)
When a function is made of compositions and concatenations of simple functions, those two rules enable efficient computation of the Gramian. In particular, let ${\boldsymbol{f}}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{\ell}\to{\mathbb{R}}^{m}$ be some layer of a sequential neural network. Given an input ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, some parameter ${\boldsymbol{p}}\in{\mathbb{R}}^{\ell}$, the output is given by ${\boldsymbol{y}}={\boldsymbol{f}}({\boldsymbol{x}},{\boldsymbol{p}})$. If $G\in\mathcal{M}_{n}$ is some Gramian associated to ${\boldsymbol{x}}$, then forwarding $G$ through ${\boldsymbol{f}}$ yields
$\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}({\boldsymbol{x}},{\boldsymbol{p}})\cdot G\cdot\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}({\boldsymbol{x}},{\boldsymbol{p}})^{\top}+\mathcal{G}\kern-1.0pt_{2}{\boldsymbol{f}}({\boldsymbol{x}},{\boldsymbol{p}}).$ (12)
Appendix F contains examples of Gramian forwarding for some usual layers. It is worth noting that the matrices of (12) can be very sparse and structured. For this reason, representing them as arrays can be inefficient. Finding the optimal representations in all practical cases and devising a full algorithm for an optimized computation of the Gramian remains an open challenge.
Still, these considerations have the potential to drastically reduce the computation time of JD and SSJD for some architectures. This would make our methods more extensively applicable and would thus be a significant step towards unlocking multi-objective optimization as a practical paradigm.

### v2

@@ 4 Existing aggregators
In the context of multi-task learning, several works have proposed iterative optimization algorithms based on the combination of task-specific gradients (Sener & Koltun, 2018; Yu et al., 2020; Liu et al., 2021b, a; Lin et al., 2021; Navon et al., 2022; Senushkin et al., 2023). These methods can be formulated as variants of JD parameterized by different aggregators. More specifically, since the gradients are stochastically estimated from batches of data, these are cases of what we call SEJD. In the following, we briefly present the most prominent aggregators and summarize their properties in Table 1. As a baseline, we also consider $\mathcal{A}_{\text{Mean}}$, which simply averages the rows of the Jacobian. Their formal definitions are provided in Appendix B. Some of them are also illustrated in Figure 1(b).
@@
$\mathcal{A}_{\text{RGW}}$ aggregates the matrix using a random vector of weights (Lin et al., 2021). $\mathcal{A}_{\text{MGDA}}$ gives the aggregation that maximizes the smallest improvement (Désidéri, 2012; Sener & Koltun, 2018; Fliege & Svaiter, 2000). $\mathcal{A}_{\text{CAGrad}}$ maximizes the smallest improvement in a ball around the average gradient whose radius is parameterized by $c\in\left[0,1\right[{}$ (Liu et al., 2021a). $\mathcal{A}_{\text{PCGrad}}$ projects each gradient onto the orthogonal hyperplane of other gradients in case of conflict, iteratively and in a random order (Yu et al., 2020). It is, however, only non-conflicting when $m\varleq 2$, in which case $\mathcal{A}_{\text{PCGrad}}=m\cdot\mathcal{A}_{\text{UPGrad}}$. IMTL-G is a method to balance some gradients with impartiality (Liu et al., 2021b). It is only defined for linearly independent gradients, but we generalize it as a formal aggregator, denoted $\mathcal{A}_{\text{IMTL-G}}$, in Appendix B.6. Aligned-MTL orthonormalizes the Jacobian and weights its rows according to some preferences (Senushkin et al., 2023). We denote by $\mathcal{A}_{\text{Aligned-MTL}}$ this method with uniform preferences. $\mathcal{A}_{\text{Nash-MTL}}$ aggregates Jacobians by finding the Nash equilibrium between task-specific gradients (Navon et al., 2022). Lastly, the GradDrop layer (Chen et al., 2020) defines a custom backward pass that combines gradients with respect to some internal activation. The corresponding aggregator, denoted $\mathcal{A}_{\text{GradDrop}}$, randomly drops out some gradient coordinates based on their sign and sums the remaining ones.
@@
In the context of continual learning, to limit forgetting, an idea is to project the gradient onto the dual cone of gradients computed with past examples (Lopez-Paz & Ranzato, 2017). This idea can be translated into an aggregator that projects the mean gradient onto the dual cone of the rows of the Jacobian. We name this $\mathcal{A}_{\text{DualProj}}$.
@@
Several other works consider the gradients to be noisy when making their theoretical analysis (Liu & Vicente, 2021; Zhou et al., 2022; Fernando et al., 2022; Chen et al., 2024; Xiao et al., 2024). Their solutions for combining gradients are typically stateful. Unifying their formulations would thus require a more complex variant of Jacobian descent.
@@
In the federated learning setting, several aggregators have been proposed to combine the model updates while being robust to adversaries (Blanchard et al., 2017; Guerraoui et al., 2018; Chen et al., 2017; Yin et al., 2018). We do not study them here as they mainly focus on security aspects.
@@
Désidéri (2012) $\mathcal{A}_{\text{MGDA}}$
Lopez-Paz & Ranzato (2017) $\mathcal{A}_{\text{DualProj}}$
Yu et al. (2020) $\mathcal{A}_{\text{PCGrad}}$
Chen et al. (2020) $\mathcal{A}_{\text{GradDrop}}$
Liu et al. (2021b) $\mathcal{A}_{\text{IMTL-G}}$
Liu et al. (2021a) $\mathcal{A}_{\text{CAGrad}}$
Lin et al. (2021) $\mathcal{A}_{\text{RGW}}$
Navon et al. (2022) $\mathcal{A}_{\text{Nash-MTL}}$
Senushkin et al. (2023) $\mathcal{A}_{\text{Aligned-MTL}}$
@@ 5 Experiments
In the following, we present empirical results for instance-wise risk minimization on some simple image classification datasets. IWRM is performed by stochastic sub-Jacobian descent, as described in Section 3. A key consideration is that when the aggregator is $\mathcal{A}_{\text{Mean}}$, this approach becomes equivalent to empirical risk minimization with SGD. It is thus used as a baseline for comparison.
We train convolutional neural networks on subsets of SVHN (Netzer et al., 2011), CIFAR-10 (Krizhevsky et al., 2009), EuroSAT (Helber et al., 2019), MNIST (LeCun et al., 1998), Fashion-MNIST (Xiao et al., 2017) and Kuzushiji-MNIST (Clanuwat et al., 2018). To make the comparisons as fair as possible, we have tuned the learning rate very precisely for each aggregator, as explained in detail in Appendix C.1. We have also run the same experiments several times independently to gain confidence in our results. Since this leads to a total of $43776$ training runs across all of our experiments, we have limited the size of each training dataset to 1024 images, greatly reducing computational costs. Note that this is strictly an optimization problem: we are not studying the generalization of the model, which would be captured by some performance metric on a test set. Other experimental settings, such as the network architectures and the total computational budget used to run our experiments, are given in Appendix C. Figure 2 reports the main results on SVHN and CIFAR-10, two of the datasets exhibiting the most substantial performance gap. Results on the other datasets and aggregators are reported in Appendix D.1. They also demonstrate a significant performance gap.
@@
Figure 2: Optimization metrics obtained with IWRM with 1024 training examples and a batch size of 32, averaged over 8 independent runs. The shaded area around each curve shows the estimated standard error of the mean over the 8 runs. Curves are smoothed for readability. Best viewed in color.
Here, we compare the aggregators in terms of their average loss over the training set: the goal of ERM. For this reason, it is rather surprising that $\mathcal{A}_{\text{Mean}}$, which directly optimizes this objective, exhibits a slower convergence rate than some other aggregators. In particular, $\mathcal{A}_{\text{UPGrad}}$, and to a lesser extent $\mathcal{A}_{\text{DualProj}}$, provide improvements on all datasets.
Figures 2(b) and 2(d) show the similarity between the update of each aggregator and the update given by $\mathcal{A}_{\text{Mean}}$. For $\mathcal{A}_{\text{UPGrad}}$, a low similarity indicates that there are some conflicting gradients with imbalanced norms (a setting illustrated in Figure 1). Our interpretation is thus that $\mathcal{A}_{\text{UPGrad}}$ prevents gradients of hard examples from being dominated by those of easier examples early into the training. Since fitting those is more complex and time-consuming, it is beneficial to consider them earlier. We believe the similarity increases later on because the gradients become more balanced. This further suggests a greater stability of $\mathcal{A}_{\text{UPGrad}}$ compared to $\mathcal{A}_{\text{Mean}}$, which may allow it to perform effectively at a higher learning rate and, consequently, accelerate its convergence.
The sub-optimal performance of $\mathcal{A}_{\text{MGDA}}$ in this setting can be attributed to its sensitivity to small gradients. If any row of the Jacobian approaches zero, the aggregation by $\mathcal{A}_{\text{MGDA}}$ will also approach zero. This observation illustrates the discrepancy between stationarity and optimality, as discussed in Section 2.4. A notable advantage of linearity under scaling is to explicitly prevent this from happening.
Overall, these experiments demonstrate a high potential for the IWRM paradigm and confirm the relevance of JD, and more specifically of SSJD, as multi-objective optimization algorithms. Besides, the superiority of $\mathcal{A}_{\text{UPGrad}}$ in such a simple setting supports our theoretical results.
While increasing the batch size in SGD reduces variance, the effect of doing so in SSJD combined with $\mathcal{A}_{\text{UPGrad}}$ is non-trivial, as it also tightens the dual cone. Additional results obtained when varying the batch size or updating the parameters with the Adam optimizer are available in Appendices D.2 and D.3, respectively.
While an iteration of SSJD is more expensive than an iteration of SGD, its runtime is influenced by several factors, including the choice of aggregator, the parallelization capabilities of the hardware used for Jacobian computation, and the implementation. Appendix E provides an empirical comparison of the runtime of SGD and SSJD at the current state of our library, for various aggregators. Additionally, we propose a path towards a more efficient implementation in the next section.
6 Gramian-based Jacobian descent
@@
When the number of objectives is dominated by the number of parameters of the model, the main overhead of JD comes from the usage of a Jacobian matrix rather than a single gradient. In the following, we motivate an alternative implementation of JD that only uses the inner products between each pair of gradients.
@@
For any $J\in\mathbb{R}^{m\times n}$, the matrix $G=JJ^{\top}$ is called the Gramian of $J$ and is positive semi-definite. Let $\mathcal{M}_{m}\subseteq\mathbb{R}^{m\times m}$ be the set of positive semi-definite matrices. The Gramian of the Jacobian, denoted $\mathcal{G}\kern-1.0pt{\bm{f}}({\bm{x}})=\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})^{\top}\in\mathcal{M}_{m}$, captures the relations – including conflicts – between all pairs of gradients. Whenever $\mathcal{A}$ is a weighted aggregator, the update of JD is ${\bm{y}}=-\eta\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})^{\top}{\bm{w}}$ for some vector of weights ${\bm{w}}\in{\mathbb{R}}^{m}$. Substituting this into the Taylor approximation of (2) gives
${\bm{f}}({\bm{x}}+{\bm{y}})={\bm{f}}({\bm{x}})-\eta\mathcal{G}\kern-1.0pt{\bm{f}}({\bm{x}})\cdot{\bm{w}}+o\left(\eta\sqrt{{\bm{w}}^{\top}\cdot\mathcal{G}\kern-1.0pt{\bm{f}}({\bm{x}})\cdot{\bm{w}}}\right).$ (9)
@@
This expression only depends on the Jacobian through its Gramian. It is thus sensible to focus on aggregators whose weights are only a function of the Gramian. Denoting this function as $\mathcal{W}~{}:~{}\mathcal{M}_{m}\to{\mathbb{R}}^{m}$, those aggregators satisfy $\mathcal{A}(J)=J^{\top}\cdot\mathcal{W}(G)$. Remarkably, all weighted aggregators of Table 1 can be expressed in this form. In the case of $\mathcal{A}_{\text{UPGrad}}$, this is clearly demonstrated in Proposition 1, which shows that the weights depend on $G$. For such aggregators, substitution and linearity of differentiation [^3] For any ${\bm{x}}\in{\mathbb{R}}^{n}$ and any ${\bm{w}}\in{\mathbb{R}}^{m}$, $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})^{\top}{\bm{w}}=\nabla\left({\bm{w}}^{\top}{\bm{f}}\right)({\bm{x}})$ then yield
$\mathcal{A}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\big{)}=\nabla\Big{(}\mathcal{W}\big{(}\mathcal{G}\kern-1.0pt{\bm{f}}({\bm{x}})\big{)}^{\top}\cdot{\bm{f}}\Big{)}({\bm{x}}).$ (10)
@@
After computing $\mathcal{W}\big{(}\mathcal{G}\kern-1.0pt{\bm{f}}({\bm{x}})\big{)}$, a step of JD would thus only require the backpropagation of a scalar function. The computational cost of applying $\mathcal{W}$ depends on the aggregator and is often dominated by the cost of computing the Gramian.
We now outline a method to compute the Gramian of the Jacobian without ever having to store the full Jacobian in memory. Similarly to the backpropagation algorithm, we can leverage the chain rule. Let ${\bm{g}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{k}$ and ${\bm{f}}:{\mathbb{R}}^{k}\to{\mathbb{R}}^{m}$, then for any ${\bm{x}}\in{\mathbb{R}}^{n}$, the chain rule for Gramians is
$\mathcal{G}\kern-1.0pt({\bm{f}}\circ{\bm{g}})({\bm{x}})=\mathcal{J}\kern-2.0pt{\bm{f}}\big{(}{\bm{g}}({\bm{x}})\big{)}\cdot\mathcal{G}\kern-1.0pt{\bm{g}}({\bm{x}})\cdot\mathcal{J}\kern-2.0pt{\bm{f}}\big{(}{\bm{g}}({\bm{x}})\big{)}^{\top}.$ (11)
@@
Moreover, when the function has multiple inputs, the Gramian can be computed as a sum of individual Gramians. Let ${\bm{f}}:\mathbb{R}^{n_{1}+\cdots+n_{k}}\to{\mathbb{R}}^{m}$ and ${\bm{x}}=\begin{bmatrix}{\bm{x}}_{1}^{\top}&\cdots&{\bm{x}}_{k}^{\top}\end{bmatrix}^{\top}$. We can write $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})$ as the concatenation of Jacobians $\begin{bmatrix}\mathcal{J}\kern-2.0pt_{{\bm{x}}_{1}}{\bm{f}}({\bm{x}})&\cdots&\mathcal{J}\kern-2.0pt_{{\bm{x}}_{k}}{\bm{f}}({\bm{x}})\end{bmatrix}$, where $\mathcal{J}\kern-2.0pt_{{\bm{x}}_{i}}{\bm{f}}({\bm{x}})$ is the Jacobian of ${\bm{f}}$ with respect to ${\bm{x}}_{i}$ evaluated at ${\bm{x}}$. For any $i\in[k]$, let $\mathcal{G}\kern-1.0pt_{{\bm{x}}_{i}}{\bm{f}}({\bm{x}})=\mathcal{J}\kern-2.0pt_{{\bm{x}}_{i}}{\bm{f}}({\bm{x}})\cdot\mathcal{J}\kern-2.0pt_{{\bm{x}}_{i}}{\bm{f}}({\bm{x}})^{\top}$. Then
$\mathcal{G}\kern-1.0pt{\bm{f}}({\bm{x}}_{1},\dots,{\bm{x}}_{k})=\sum_{i\in[k]}\mathcal{G}\kern-1.0pt_{{\bm{x}}_{i}}{\bm{f}}({\bm{x}}_{1},\dots,{\bm{x}}_{k}).$ (12)
When a function is made of compositions and concatenations of elementary functions, the Gramian of the Jacobian can thus be expressed with sums and products of partial Jacobians.
We now provide an example algorithm to compute the Gramian of a sequence of layers. For $0\varleq i<k$, let ${\bm{f}}_{i}:{\mathbb{R}}^{n_{i}}\times{\mathbb{R}}^{\ell_{i}}\to{\mathbb{R}}^{n_{i+1}}$ be a layer parameterized by ${\bm{p}}_{i}\in{\mathbb{R}}^{\ell_{i}}$. Given ${\bm{x}}_{0}\in{\mathbb{R}}^{n_{0}}$, for $0\varleq i<k$, the activations are recursively defined as ${\bm{x}}_{i+1}={\bm{f}}_{i}({\bm{x}}_{i},{\bm{p}}_{i})$. Algorithm 3 illustrates how (11) and (12) can be combined to compute the Gramian of the network with respect to its parameters.
$J_{x}\leftarrow I$    # Identity matrix of size $n_{k}\times n_{k}$
$G\hskip 2.56064pt\leftarrow 0$   # Zero matrix of size $n_{k}\times n_{k}$
for $i\leftarrow k-1$ to 00 do
       $J_{p}\leftarrow\mathcal{J}\kern-2.0pt_{{\bm{p}}_{i}}{\bm{f}}_{i}({\bm{x}}_{i},{\bm{p}}_{i})\cdot J_{x}$   # Jacobian of ${\bm{x}}_{k}$ w.r.t. ${\bm{p}}_{i}$
       $J_{x}\leftarrow\mathcal{J}\kern-2.0pt_{{\bm{x}}_{i}}{\bm{f}}_{i}({\bm{x}}_{i},{\bm{p}}_{i})\cdot J_{x}$   # Jacobian of ${\bm{x}}_{k}$ w.r.t. ${\bm{x}}_{i}$
       $G\hskip 2.56064pt\leftarrow G+J_{p}J_{p}^{\top}$
Output: $G$
Algorithm 3 Gramian reverse accumulation for a sequence of layers
Generalizing Algorithm 3 to any computational graph and implementing it efficiently remains an open challenge extending beyond the scope of this work.